Find the sum of:
(a) 137 and -354
(b) -52 and 52 .
(d) -312, 39 and 192
(d) -50, -200 and 300

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Solution: (a) 137 and -354 (137) + ( -354) = (137) + (-137) + (-217) [∵ (137) + (-137) = 0] = 0 + (-217) = (-217) (b) -52 and 52 (-52) + (+52) = 0 [ ∵ (-a) + (+ a) = 0] (c) -312, 39 and 192 (-312) + (+ 39) + (+192) = (-231) + (- 81) + (+39) + (+192) = (-231) + (-81) + (+ 231) = (-231) + (+ 231) + (-81) [∵ (-a) + (a) = 0] = 0 + (- 81) = – 81 (d) – 50, -200 and 300 (-50) + (-200) + (+ 300) = (-50) + (-200) + (+200) + (+100) = (-50) + 0 + (+100)[∵ (-a) + (+ a) = 0] = (-50) + (+100) = (-50) + (+50) + (+ 50) = 0 + (+ 50) = 50 [ ∵ (-a) + (+ a) = 0]